OpenStudy (anonymous):

implicit differentiation for: (2x-3x)^2=x^2+y^2? help...

OpenStudy (anonymous):

Differentiate each term. Differentiate y and tag on a dy/dx as a chain rule. Solve for dy/dx

OpenStudy (anonymous):

assuming \[(2x-3y)^2 =x^2 +y^2\]

OpenStudy (anonymous):

differentiate both sides with respect to x

OpenStudy (anonymous):

\[2(2x-3y) ( 2 -3\frac{dy}{dx}) =2x +2y \frac{dy}{dx}\]

OpenStudy (anonymous):

the term on the LHS had chain rule applied to it now just solve for dy/dx you can do that, its just algebra